OpenStudy (anonymous):

The sum of ten integers is 32. What is the smallest of the ten? A) -34 B) -32 C) -10 D) -2 E) It cannot be determined from the information given

OpenStudy (anonymous):

there isnt enough information in this question, right? the problem doesnt specify enough what the numbers have to be...

OpenStudy (anonymous):

im sorry, im slightly confused... how does x1 = x2+1?

OpenStudy (anonymous):

32=x1+ x2 +x3+ x4+x5+x6+x7+x8+x9+x10, x2=x1+1, x3=x2+1, x4=x3+1, x5=x4+1, x6=x5+1, x7=x6+1, x8=x7+1, x9=x8+1, x10=x9+1,

OpenStudy (anonymous):

hold on imma check if i get an answer then i'll explain my logic

OpenStudy (anonymous):

yeah i saw the pattern you were trying to get at but... i would really appreciate some explanation XD i think i missed this one on the last practice SAT

OpenStudy (anonymous):

thanks for being helpful btw

OpenStudy (anonymous):

32=x1+ x2 +x3+ x4+x5+x6+x7+x8+x9+x10 well this part is because you have 10 numbers that add up to 32

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so then each one is 1plus more than the least except for the first one which we will choose as x1 so there fore x2=x1+1 and since the next one x3 and it has to be plus 1 more than the last it is x3=x2+1 and so on and so on? right?

OpenStudy (anonymous):

Oh ok so thats why you added 1. Because its the least number... yeah that makes much more sense

OpenStudy (anonymous):

i get -1.30 it is odd since it is not a choice

OpenStudy (anonymous):

This question is ridiculous... ah well, thanks for being persistant in helping to get an answer

OpenStudy (anonymous):

it is -1.3 try it out add consecutive numbers to it and see if you get 36. The otther thing to do is get A) -34 B) -32 C) -10 D) -2 and plug those in as well to see if you get 36..

OpenStudy (anonymous):

E it could be e cuzz i was thinking consecutive numbers but it doesn't say consecutive anywhere but i feel it is implying it.

OpenStudy (anonymous):

Right, thats what i originally thought from the start hahahaha Again, thanks a lot! :)

OpenStudy (anonymous):

Timo is wrong, and you were right. There is not enough information because it does not say that they are consecutive integers.

OpenStudy (anonymous):

that did occur to me look above :P